It has been over 40 years since I have had algebra. Below is a problem my son needs to solve. If you can provide the answers hopefully I will begin to remember how equations work so I can help him with the rest of his problems. Thanks.
The equation for profit P in terms of the number of boxes sold s is:
P = 5s - (100 + 2s). The number of boxes sold also depends on the number of choir members.
a. Suppose each member sell 11 boxes. Write an equation that will predict profit from the number of choir members n. (Hint: Firts find an expression for the number of boxes sold)
b. Write an equivalent expression for profit in part (a). Explain what the variables and numbers represent.
c. Suppose the choir has 47 members. What is the profit?
d. Suppose the profit is $1,088. How many choir members are there?
e. In part (d), how many boxes were sold?
Barbara Best

Found 2 solutions by richwmiller, scott8148:
Answer by richwmiller(17219)

I suspect that there is something wrong with your equation/formula. It can be reduced. Formulas are in their reduced form
P=5s-(100+2s)
P=3s-100

Answer by scott8148(6628)

a. n members selling 11 boxes each, so 11n boxes sold ___ s = 11n ___ P = 5(11n) - [100 + 2(11n)]

b. P = 55n - (100 + 22n)
___ P is profit
___ 55 is amount collected by each choir member
___ n is number of choir members
___ 100 is fixed costs of program (marketing fees?)
___ 22 is amount paid to supplier (marketer) from each choir member

c. P = 33(47) - 100

d. 1088 = 33n - 100 ___ 1188 = 33n ___ 1188/33 = n

e. s still equals 11n, so 11 times answer d.